Mathematics 243 – Algebraic Structures Definitions Quiz 7 – October 27, 2006

Your Name:

A) (5) What is the definition of the relation x ≡ y mod n for integers x, y, n?

B) (5) Let f be the mapping that takes an integer x to its congruence class mod 23: f(x) = [x] for the congruence mod 23 relation. Is f a one-to-one mapping? Is f an onto mapping to the set of all congruence classes mod 23?